1. A sample of 25.0 mL of 0.120 M Ba(OH)2 is titrated with 0.150 M HBr. What volume of hydrobromic acid is needed to completely neutralize barium hydroxide?
1
Expert's answer
2020-12-06T13:52:16-0500

V(Ba(OH)2)=25.0mL=0.0250L

n(Ba(OH)2)= c*V = 0.0250L * 0.120mol/L = 0.003mol


Ba(OH)2 + 2HBr = BaBr2 + 2H2O


n(HBr)=2n(Ba(OH)2)=0.006mol

V(HBr)=n/c=0.006mol / 0.150mol/L = 0.040L = 40 mL